OpenStudy (anonymous):

Find d if 2(d - 4) = -2d + 8 F. -4 G. -2 H. 2 I. 4

OpenStudy (anonymous):

2(d - 4) = - 2d + 8 (multiply 2 through the parenthesis) 2d - 8 = - 2d + 8 2d + 2d = 8 + 8 4d = 16 d = 4 check... 2(d - 4) = - 2d + 8 2(4 - 4) = - 2(4) + 8 2(0) = -8 + 8 0 = 0 (correct) ANSWER : d = 4

OpenStudy (anonymous):

thats is the same thing that i got thank you for verigying that u deserve a medal

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

your welcome if u continue helping me with accurate details n answers u will have lots of medals